# Integrated Sensing and Edge AI: Realizing Intelligent Perception in 6G

### v1

@@ Abstract
Sensing and edge artificial intelligence (AI) are envisioned as two essential and interconnected functions in sixth-generation (6G) mobile networks. On the one hand, sensing-empowered applications rely on powerful AI models to extract features and understand semantics from ubiquitous wireless sensors. On the other hand, the massive amount of sensory data serves as the fuel to continuously refine edge AI models. This deep integration of sensing and edge AI has given rise to a new task-oriented paradigm known as integrated sensing and edge AI (ISEA), which features a holistic design approach to communication, AI computation, and sensing for optimal sensing-task performance. In this article, we present a comprehensive survey for ISEA. We first provide technical preliminaries for sensing, edge AI, and new communication paradigms in ISEA. Then, we study several use cases of ISEA to demonstrate its practical relevance and introduce current standardization and industrial progress. Next, the design principles, metrics, tradeoffs, and architectures of ISEA are established, followed by a thorough overview of ISEA techniques, including digital air interface, over-the-air computation, and advanced signal processing. Its interplay with various 6G advancements, e.g., new physical-layer and networking techniques, are presented. Finally, we present future research opportunities in ISEA, including the integration of foundation models, convergence of ISEA and integrated sensing and communications (ISAC), and ultra-low-latency ISEA.
@@ I Introduction
Although the coexistence and interdependence of communications, AI, and sensing is a natural trend in 6G, their cooperation and integrated design for real-timeness and high reliability (e.g., 30 ms inference latency and near 100% accuracy for autonomous driving[10]) is faced with several critical challenges, including communication and computation capacity, operational complexity, security and privacy, and sustainability. For instance, it is essential for current networking architectures and air interface techniques to evolve to accommodate the demanding requirements in latency and reliability of transmission and analytics of sensory data. These challenges necessitate innovative approaches for resource management, integrated frameworks, signal processing techniques, networking architectural design, etc., all unified under a new set of metrics and principles oriented towards optimized performance of sensing applications. This gives rise to a new task-oriented paradigm, integrated sensing and edge AI (ISEA), which is the theme of this paper. In what follows, we first provide a historical perspective of ISEA, presenting the milestones of communications, sensing, and AI to elaborate on the grand convergence of these three domains, and demonstrate how 6G, as a unified communication-computing platform, can support integrated sensing and AI in a harmonious manner. After that, we present the major contributions of this paper with comparisons to other works and the paper’s organization.
@@ I-B Evolving towards Integrated Sensing and Edge AI
Figure 1: Differences and overlappings in terms of scopes of ISEA compared with ISAC and SemCom. Compared with ISAC, which focuses on the integration and mutual assistance of RF sensing and communication, ISEA features multiple sensing modalities and joint communication-computation design for optimized performance of AI-based sensing tasks, serving applications that require rich environmental semantics. Compared with SemCom, ISEA targets specific sensing tasks and involves the joint design of a broader range of operations and components.
@@ I-C Comparisons with Relevant Works and Our Contribution
  • Comparisons with ISAC: ISAC jointly designs RF sensing and communication, improving resource utilization through their coexistence and mutual assistance within the same hardware, spectrum and waveform[18, 19]. Both ISEA and ISAC view sensing as a built-in function of the mobile network, not an independent module built on separate devices that interface with the network’s application layer. Also, both paradigms exploit radio-frequency (RF) sensing potentially with multi-functional waveforms in millimeter wave (mmWave) or terahertz (THz) frequency bands. Positioning, ranging and tracking are viewed as important application scenarios by both ISEA and ISAC.
@@
  • Comparisons with SemCom: SemCom goes beyond the traditional paradigm of building reliable bit pipes, aiming instead at the efficient transmission of information semantics via AI-based encoder-decoders and shared knowledge bases[20, 21]. The greatest common point between ISEA and SemCom is the native integration of AI models into the communication system to understand, compress and reconstruct semantics from multi-modal data for efficient communication beyond Shannon’s limit.
@@
Several surveys and reviews are related to the topic of ISEA [9, 23, 24, 5, 25, 19, 26]. [26] introduces the integration of AI, IoT, and edge computing, surveys its potential applications, and enables techniques for AI inference in IoT with end-edge-cloud orchestration as well as decentralized edge learning. In [5], the authors review energy-efficient design techniques for edge AI operations, i.e., data acquisition, training and inference, to fit AI capabilities into resource-limited wireless edge networks. However, theses works [26, 5] focus on the general edge AI where sensing is one of its applications without a thorough discussion on sensing-oriented techniques. From the sensing perspective, [25] investigates the design issues for collaborative sensing including sensing models, deployment and scheduling techniques, and metrics. However, the main focus is solutions for sensing coverage maximization, not covering wireless link issues or downstream inference. [19] discusses the advances in joint radar sensing and communications with applications to localization and sensing applications in IoT, which, however, does not involve other sensing modalities or AI-empowered data processing. The integration between RF sensing, multi-modal sensing and communication modules, named Synesthesia of Machines (SoM), is introduced in [24], which reviews current progress on the mapping between sensory data and channel data as well as mutual assistance of sensing and communications. Nevertheless, its main focus is the relationship and interaction between RF and other sensing modalities in the coexistence of communicational functions, not a generalized framework for efficient sensory data processing and exchange as in ISEA. The integration of sensing, communication and computation (ISCC), as reviewed in [23, 9], is an emerging framework exploiting coupling of the three operations for a wide range of performance metrics, e.g., 6G communication rates and E2E task latency. Compared with [23, 9], our paper focuses on the perspective on the native integration of AI models into sensing in edge networks oriented towards the performance of AI-empowered sensing tasks. Our main contributions are summarized as follows.
@@ I-D Paper Organization
The remainder of this paper is organized as follows. We first introduce useful preliminary knowledge for ISEA involving sensing data acquisition, edge AI and new communication technologies, respectively in Section II-A, Section II-B and Section II-C. Then, in Section III, we provide case studies for ISEA ranging from its application scenarios and supporting techniques that fit into the ISEA paradigm, introduce the standardization efforts for and industrial perspectives on ISEA, and available datasets. The key design principles of ISEA are defined in Section IV-A, followed by the metrics and tradeoffs in ISEA in Section IV-B. The general architectures of ISEA are introduced in Section V-A, and several specific ISEA paradigms are sketched in Section V-B. ISEA techniques based on digital air interface are presented in Section VI, while those based on AirComp air interfaces are presented in Section VII. The incorporation of advanced signal processing techniques into ISEA is introduced in Section VIII. We discuss opportunities to boost ISEA using other trending 6G advancements and review relevant works in Section IX. Outlooks for research opportunities and open issues are provided in Section X. Finally, we conclude this paper in Section XI. For the readers’ convenience, we have included Fig. 2 to illustrate the outline of this paper.
@@ II-A5 Multi-modal sensing
A rich literature exists on sensing models that exploit multiple sensing modalities to complete a downstream task. The fusion of LiDAR and camera image is arguably the most prevalent fusion strategy applied for various tasks including object detection, tracking, and semantic segmentation [45, 46, 47, 48]. The authors of[49] complement LiDAR with RF sensing to improve the detection performance. Depending on the model architecture design, the fusion methods can happen at different levels, e.g., raw-data level, feature level, and decision level. Generally, modality fusion at higher levels results in a smaller communication volume for information exchange between sensors, while the information loss may increase due to the reduced data dimension.
@@ II-B2 Edge Inference
Whilst edge learning aims at distilling intelligence from data collected by edge devices, edge inference focuses on the efficient provision of well-trained AI models as services to heterogeneous edge devices under possibly stringent latency and reliability requirements. In contrast to the conventional cloud inference which places the intensive computation load of AI model inference at the central cloud, edge inference deploys computation resources on edge nodes in proximity to the device requesting service, thus promising lower round-trip latency and avoiding traffic jams in the backbone network. These advantages render edge inference as a key enabling technology for sensing with latency-demanding and mission-critical natures. Depending on the deployment of the model computation load, edge inference can be categorized into on-device inference, on-server inference, split inference, and distributed inference. On-device and on-server inference places model computation solely on end devices or servers, aiming at scenarios with extremely scarce communication resources or low-end on-device computation capabilities, respectively. Split inference is arguably the most prevalent approach where the AI model is split into a low-complexity device sub-model and a server sub-model due to its privacy-preserving nature and high flexibility. The device uses its sub-model to extract features from the raw data, which are uploaded to the server for further inference with the server sub-model. Therein, the feature uploading phase is often confronted with a communication bottleneck since the extracted features may have a higher dimension than raw data, especially when handling bursting arrivals of inference requests from multiple users. A set of techniques have been developed for split inference to boost its efficiency under QoS requirements, including communication and computation resource allocation[56, 57], feature compression/pruning[58, 59], JSCC [60], progressive transmission[61], early exiting[62, 63], and model partitioning point selection[64]. Different from previous paradigms involving a device-server pair, distributed inference involves the distribution of computation load onto multiple edge devices to complete a single inference request [65, 66], which usually requires the partition of an AI model into multiple independent execution units for distributed computing.
@@ III-B4 Human-Machine Symbiosis
Advanced sensor technologies, such as IoT devices, wearable sensors, and smart environments, enable a seamless flow of data between humans and machines. These techniques, coupled with AI-based data analytics, offer personalized and responsive interactions for cooperative interaction and close coupling between humans and machines [21], known as human-machine symbiosis. Several practical examples illustrate how AI-sensing integrated techniques are intricately woven into various facets of our lives, enriching human experiences and decision-making. For instance, leveraging sensory data (e.g., based on motion and temperature sensors, thermostats, cameras, or wireless sensing) within smart homes enables automation based on the recognition of human behaviour (e.g., voice and gestures), adjusting the environment according to occupants’ preferences and optimizing energy efficiency and comfort [21]. ISEA, propelled by and designed for low-latency intelligent sensory data analytics, can serve as a powerful platform for collectively implementing sensing techniques and AI to offer real-time human-machine symbiosis services.
III-C Standardization Efforts, Industrial Perspective, and Datasets
III-C1 Standardization Efforts

### v2

@@ Abstract
Sensing and edge artificial intelligence (AI) are envisioned as two essential and interconnected functions in sixth-generation (6G) mobile networks. On the one hand, sensing-empowered applications rely on powerful AI models to extract features and understand semantics from ubiquitous wireless sensors. On the other hand, the massive amount of sensory data serves as the fuel to continuously refine edge AI models. This deep integration of sensing and edge AI has given rise to a new task-oriented paradigm known as integrated sensing and edge AI (ISEA), which features a holistic design approach to communication, AI computation, and sensing for optimal sensing-task performance. In this article, we present a comprehensive survey for ISEA. We first provide technical preliminaries for sensing, edge AI, and new communication paradigms in ISEA. Then, we study several use cases of ISEA to demonstrate its practical relevance and introduce current standardization and industrial progress. Next, the design principles, metrics, tradeoffs, and architectures of ISEA are established, followed by a thorough overview of ISEA techniques, including digital air interface, over-the-air computation, and advanced signal processing. Its interplay with various 6G advancements, e.g., new physical-layer and networking techniques, are presented. Finally, we present future research opportunities in ISEA, including the integration of foundation models, convergence of ISEA and integrated sensing and communications (ISAC), ultra-low-latency ISEA, and practicality issues.
@@ I Introduction
Although the coexistence and interdependence of communications, AI, and sensing is a natural trend in 6G, their cooperation and integrated design for real-timeness and high reliability (e.g., 30 ms inference latency and near 100% accuracy for autonomous driving[10, 11]) is faced with several critical challenges, including communication and computation capacity, operational complexity, security and privacy, and sustainability. For instance, it is essential for current networking architectures and air interface techniques to evolve to accommodate the demanding requirements in latency and reliability of transmission and analytics of sensory data. These challenges necessitate innovative approaches for resource management, integrated frameworks, signal processing techniques, networking architectural design, etc., all unified under a new set of metrics and principles oriented towards optimized performance of sensing applications. This gives rise to a new task-oriented paradigm, integrated sensing and edge AI (ISEA), which is the theme of this paper. In what follows, we first provide a historical perspective of ISEA, presenting the milestones of communications, sensing, and AI to elaborate on the grand convergence of these three domains, and demonstrate how 6G, as a unified communication-computing platform, can support integrated sensing and AI in a harmonious manner. After that, we present the major contributions of this paper with comparisons to other works and the paper’s organization.
@@ I-B Evolving towards Integrated Sensing and Edge AI
Figure 1: Differences and overlappings in terms of the scopes of ISEA compared with ISAC and SemCom. Compared with ISAC, which focuses on the integration and mutual assistance of RF sensing and communication, ISEA features multiple sensing modalities and joint communication-computation design for optimized performance of AI-based sensing tasks, serving applications that require rich environmental semantics. Compared with SemCom, ISEA targets specific sensing tasks and involves the joint design of a broader range of operations and components.
@@ I-C Comparisons with Relevant Works and Our Contribution
  • Comparisons with ISAC: ISAC jointly designs RF sensing and communication, improving resource utilization through their coexistence and mutual assistance within the same hardware, spectrum and waveform[19, 20, 21]. Both ISEA and ISAC view sensing as a built-in function of the mobile network, not an independent module built on separate devices that interface with the network’s application layer. Also, both paradigms exploit radio-frequency (RF) sensing potentially with multi-functional waveforms in millimeter wave (mmWave) or terahertz (THz) frequency bands. Positioning, ranging, tracking, and satellite-air-ground network are considered as important application scenarios by both ISEA and ISAC, as depicted in Fig. 1.
@@
  • Comparisons with SemCom: SemCom goes beyond the traditional paradigm of building reliable bit pipes, aiming instead at the efficient transmission of information semantics via AI-based encoder-decoders and shared knowledge bases[22, 23]. The greatest common point between ISEA and SemCom is the native integration of AI models into the communication system to understand, compress and reconstruct semantics from multi-modal data for efficient communication beyond Shannon’s limit. As illustrated in Fig. 1, both paradigms are empowered by a common set of enabling techniques, including wireless architectures for distributed inference and learning, semantic encoder-decoder frameworks, joint source-channel coding, and the adoption of foundation models and generative AI for enhanced semantic representation and generalization. These techniques collectively facilitate a shift from symbol-level fidelity to meaning-level accuracy, providing the platform for diverse human-machine interaction applications.
@@
Overall, the proposed ISEA framework fundamentally differs from ISAC’s resource-sharing paradigm (optimizing communication-sensing trade-offs) and SemCom’s semantic compression (balancing bandwidth against contextual distortion). By jointly optimizing AI, communication, and sensing subsystems under an end-to-end performance criterion, ISEA transforms traditional component-level trade-offs - such as data delivery precision-latency compromise or model splitting’s computation-communication balance - into tunable parameters for global optimization[25]. This system-level approach enables superior adaptability where conventional paradigms must make hard resource partitioning decisions.
TABLE I: Comparisons with Related Survey Papers
Ref. Sensing Modalities Performance Metrics Communication Techniques Design Principles Research Opportunities
[26] RF signals, LiDAR, RGB cameras, CSI-based sensing Task accuracy, latency, energy consumption AirComp, JSCC, semantic communication Task-oriented ISCC, semantic-aware optimization Semantic-aware FL, edge coordination, and wireless–AI co-design
[5] Wearables, RF sensors, RGB cameras, LiDAR, ambient sensors Energy efficiency, inference latency, accuracy Gradient compression, quantized transmission, adaptive scheduling Energy-first design, model–hardware co-optimization Generative edge AI, neuromorphic computing, adaptive resource scheduling
[27] Radar, LiDAR, RGB cameras, GPS, IMU, RF signals Beam prediction, channel estimation, sensing accuracy Sensor-aided beamforming, communication-aware sensing Cross-modal coordination, ANN-enabled fusion, data-centric design Unified SoM framework, multi-sensor fusion, dynamic dataset integration
[20] Radar, LiDAR, RF signals, ultrasonic sensors, RGB cameras Sensing accuracy, coverage, E2E latency Passive radar communication, cognitive radio, full-duplex, MIMO Functional reuse, scalable JSAC, spectrum efficiency RIS-enhanced JSAC, backscatter sensing, integrated localization
[9] Radar, LiDAR, RGB cameras, mobile crowdsensing Trade-offs among sensing, communication, and computation AirComp, ISAC, triple-functional waveforms Integrated waveform design, cross-layer orchestration ISCC in digital twins, SAGIN, and computing networks
[28] LiDAR, radar, RGB cameras, wearables, industrial and environmental sensors Latency, reliability, scalability, energy consumption Bandwidth-aware compression, lightweight IoT protocols Edge–cloud synergy, privacy-aware collaboration Cross-layer scheduling, infrastructure-free AIoT, SAGIN-based fusion
[29] RGB cameras, RF signals, WSNs, mobile sensors Coverage ratio, QoM, activation efficiency Sensor scheduling, duty cycling, cluster-based communication Coverage-aware coordination, role-based activation Probabilistic models, adaptive full-view coverage, sensor selection
Ours (ISEA) Fused multi-modal data, RGB camera, LiDAR, mmWave and UWB radar, event cameras Task-oriented metrics: perception accuracy, E2E inference latency Modality-aware JSCC, AirComp, short packet transmission Task-performance co-optimization across sensing, comm, computation Integration with foundation models, low-latency ISEA, convergence of ISEA and ISAC
@@
Several surveys and reviews are related to the topic of ISEA [9, 26, 27, 5, 29, 20, 28]. Specifically, [28] introduces the integration of AI, IoT, and edge computing, surveys its potential applications, and enables techniques for AI inference in IoT with end-edge-cloud orchestration as well as decentralized edge learning. In [5], the authors review energy-efficient design techniques for edge AI operations, i.e., data acquisition, training and inference, to fit AI capabilities into resource-limited wireless edge networks. However, these works [28, 5] focus on the general edge AI where sensing is one of its applications without a thorough discussion on sensing-oriented techniques. From the sensing perspective, [29] investigates the design issues for collaborative sensing including sensing models, deployment and scheduling techniques, and metrics. However, the main focus is solutions for sensing coverage maximization, not covering wireless link issues or downstream inference. [20] discusses the advances in joint radar sensing and communications with applications to localization and sensing applications in IoT, which, however, does not involve other sensing modalities or AI-empowered data processing. The integration between RF sensing, multi-modal sensing and communication modules, named Synesthesia of Machines (SoM), is introduced in [27], which reviews current progress on the mapping between sensory data and channel data as well as mutual assistance of sensing and communications. Nevertheless, its main focus is the relationship and interaction between RF and other sensing modalities in the coexistence of communicational functions, not a generalized framework for efficient sensory data processing and exchange as in ISEA. The integration of sensing, communication and computation (ISCC), as reviewed in [26, 9], is an emerging framework exploiting coupling of the three operations for a wide range of performance metrics, e.g., 6G communication rates and E2E task latency. Compared with [26, 9], our paper focuses on the perspective on the native integration of AI models into sensing in edge networks oriented towards the performance of AI-empowered sensing tasks. Our main contributions are summarized as follows.
@@
To highlight the key distinctions of ISEA from related surveys, Table I provides a comprehensive comparison in terms of sensing modalities, performance metrics, communication techniques, design principles, and research opportunities.
@@ I-D Paper Organization
The remainder of this paper is organized as follows. We first introduce useful preliminary knowledge for ISEA involving sensing data acquisition, edge AI and new communication technologies, respectively in Section II-A, Section II-B and Section II-C. Then, in Section III, we provide case studies for ISEA ranging from its application scenarios to supporting techniques that fit into the ISEA paradigm, investigate the standardization efforts for and industrial perspectives on ISEA, and introduce available datasets. The key design principles of ISEA are defined in Section IV-A, followed by the metrics and tradeoffs in ISEA in Section IV-B. The general architectures of ISEA are introduced in Section V-A, and several specific ISEA paradigms are sketched in Section V-B. ISEA techniques based on digital air interface are presented in Section VI, while those based on AirComp air interfaces are presented in Section VII. The incorporation of advanced signal processing techniques into ISEA is introduced in Section VIII. We discuss opportunities to boost ISEA using other trending 6G advancements and review relevant works in Section IX. Outlooks for research opportunities and open issues are provided in Section X. Finally, we conclude this paper in Section XI. For the readers’ convenience, we have included Fig. 2 to illustrate the outline of this paper.
@@ II-A5 Multi-Modal Sensing
A rich literature exists on sensing models that exploit multiple sensing modalities to complete a downstream task. The fusion of LiDAR and camera image is arguably the most prevalent fusion strategy applied for various tasks including object detection, tracking, and semantic segmentation [48, 49, 50, 51]. The authors of[52] complement LiDAR with RF sensing to improve the detection performance. Depending on the model architecture design, the fusion methods can happen at different stages, e.g., early fusion at the raw-data level, intermediate fusion at the feature level, and late fusion at the decision level. Generally, modality fusion at higher levels results in a smaller communication volume for information exchange between sensors, while the information loss may increase due to the reduced data dimension. Detailed descriptions of these fusion methods and tradeoffs are shown in Fig. 4.
Figure 4: Different data fusion strategies with tradeoffs in versatility, communication overhead, computation workload, and information loss[53, 54, 55].
@@ II-B2 Edge Inference
Whilst edge learning aims at distilling intelligence from data collected by edge devices, edge inference focuses on the efficient provision of well-trained AI models as services to heterogeneous edge devices under possibly stringent latency and reliability requirements. In contrast to the conventional cloud inference which places the intensive computation load of AI model inference at the central cloud, edge inference deploys computation resources on edge nodes in proximity to the device requesting service, thus promising lower round-trip latency and avoiding traffic jams in the backbone network. These advantages render edge inference as a key enabling technology for sensing with latency-demanding and mission-critical natures. Depending on the deployment of the model computation load, edge inference can be categorized into on-device inference, on-server inference, split inference, and distributed inference. On-device and on-server inference places model computation solely on end devices or servers, aiming at scenarios with extremely scarce communication resources or low-end on-device computation capabilities, respectively. Split inference is arguably the most prevalent approach where the AI model is split into a low-complexity device sub-model and a server sub-model due to its privacy-preserving nature and high flexibility. The device uses its sub-model to extract features from the raw data, which are uploaded to the server for further inference with the server sub-model. Therein, the feature uploading phase is often confronted with a communication bottleneck since the extracted features may have a higher dimension than raw data, especially when handling bursting arrivals of inference requests from multiple users. A set of techniques have been developed for split inference to boost its efficiency under QoS requirements, including communication and computation resource allocation[62, 63], feature compression/pruning[64, 65], JSCC [66], progressive transmission[67], early exiting[68, 69], batch processing[70, 71], and model partitioning point selection[72]. Different from previous paradigms involving a device-server pair, distributed inference involves the distribution of computation load onto multiple edge devices to complete a single inference request [73, 74], which usually requires the partition of an AI model into multiple independent execution units for distributed computing.
@@ III-B4 Human-Machine Symbiosis
Advanced sensor technologies, such as IoT devices, wearable sensors, and smart environments, enable a seamless flow of data between humans and machines. These techniques, coupled with AI-based data analytics, offer personalized and responsive interactions for cooperative interaction and close coupling between humans and machines [23], known as human-machine symbiosis. Several practical examples illustrate how AI-sensing integrated techniques are intricately woven into various facets of our lives, enriching human experiences and decision-making. For instance, leveraging sensory data (e.g., based on motion and temperature sensors, thermostats, cameras, or wireless sensing) within smart homes enables automation based on the recognition of human behavior (e.g., voice and gestures), adjusting the environment according to occupants’ preferences and optimizing energy efficiency and comfort [23]. ISEA, propelled by and designed for low-latency intelligent sensory data analytics, can serve as a powerful platform for collectively implementing sensing techniques and AI to offer real-time human-machine symbiosis services.
III-C Limitations of 5G Connectivity-Centric Approaches
The conventional 5G New Radio (NR) architecture, while optimized for high-throughput communication, exhibits fundamental inefficiencies when applied to integrated sensing, AI computation, and communication. These limitations stem from three key aspects:
  • Functional isolation: 5G NR employs distinct waveform designs for communication (e.g., OFDM) and sensing (e.g., sparse reference signals), leading to: 1) Resource inefficiency: Redundant pilot transmissions for channel estimation (communication) and radar-like sensing result in extra overhead in joint positioning and tracking scenarios; 2) Latency bottlenecks: AI-driven applications (e.g., generative AI) require real-time sensor-data fusion, yet 5G’s sequential processing (sense → transmit → compute) introduces prohibitive delays for time-sensitive inference tasks.
  • Rigid resource allocation: 5G’s slot-based scheduling lacks the agility to dynamically optimize: 1) Time-frequency resources for ultra-low-latency tracking (e.g., µs-level updates in drone swarms); 2) Computing-aware transmission for distributed AI workloads (e.g., split inference across devices and edge servers).
  • Semantic-agnostic operation: Current protocols lack native support for: 1) Task-oriented QoS metrics: Traditional bit-error-rate (BER) optimizations do not align with perceptual quality in generative AI or sensing accuracy in SLAM applications; 2) Cross-domain tradeoffs: Independent optimization of communication capacity and sensing resolution (Cramér-Rao bound) leads to suboptimal performance in integrated networks.
III-D Standardization Efforts, Industrial Perspective, and Datasets
III-D1 Standardization Efforts